OpenStudy (anonymous):

Which ratio forms a proportion with 15/21

OpenStudy (anonymous):

20/26 7/5 20/28 12/18

OpenStudy (anonymous):

first reduce 15/221 we have (3*5)/ (3*7)= 5/7 now (5*4)/(7*4) = 20/28

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

welcome dear